Express as a trigonometric function of one angle.
cos 8 sin(−1) − cos 1 sin 8
Not sure how to write this. My professor only gave me double and half angle formulas so I am confused. Please help!

Answer by ikleyn(52781)

.

To solve the problem, you MUST know this formula, called "the sum of arguments for sine"

    sin(a+b) = sin(a)*cos(b) + cos(a)*sin(b)



By applying this formula, you get

    cos(8)*sin(-1) - cos(1)*sin(8) = -cos(8)*sin(1) - sin(8)*cos(1) = -(cos(8)*sin(1) + sin(8)*cos(1)) = -sin(9),


and it is your answer.


On the way, I used  sin(-1) = -sin(1).


Note that the numbers in this post (in my and in yours) are measures of angles, expressed in degrees.


Solved.

-----------------

Regarding these basic trigonometry formulas,  see the lessons
    - FORMULAS FOR TRIGONOMETRIC FUNCTIONS
    - Addition and subtraction formulas
    - Addition and subtraction formulas - Examples
in this site.

They are  REALLY  BASIC:  everybody studying  Trigonometry  MUST  know them as the multiplication table.

At least, he (or she) must know that they do exist, and also must know, where to find them . . .